ISS On-Orbit Status 1 Mar 2002


All ISS systems continue to function nominally, except as noted previously or below.

After liftoff of STS-109/Columbia at 6:22 am EST, the ISS crew took a picture with the ESC (electronic still camera), showing the vapor trail of the ascending Shuttle as seen from orbit. The image was downlinked and viewed at MCC-H only minutes after the launch.

FE-1 Carl Walz set up the IWIS (internal wireless instrumentation system), which records structural dynamics of the station. The IWIS utilizes sets of accelerometers and strain gages, both types supported by an NCU (network control unit) and their own remote sensing units (RSU) in the Lab and Node. The RSU has a small transmitter antenna to radio measurements to the SSC6 (space station computer) laptop #6 for recording on a PCMCIA hard disk card and later downlink to the ground. Today's IWIS operations were intended to record vibrations caused by crewmember exercise on the TVIS treadmill. During data take, a photo/video camera, with its field of view focused on the lower body and the entire TVIS chassis, recorded the exercise. Objective of the activity was to record the runner's footfalls, motion, and resulting TVIS displacements and frequencies. After the run, the P/TV camera setup was torn down again to be out of the way.

Reinitialization (discharge/recharČ»░▒9 of the two 4B2 batteries (4B21 and 4B22) was successfully completed. The procedure, which had been interrupted by the C&C MDM transitions yesterday, reduced the pressure difference between the two storage units to within nominal toln8ce range.

FE-1 Carl Walz worked in the Lab on the BTR (biotechnology refrigerator), repacking its contents to accommodate samples from the ADVASC mustard plant growth.

Dan Bursch supported the second attempt by the ground to download the post-EVA PuFF and EVARM data by activating and deactivating the HRF (human research facility) PC and its rack connection.

Troubleshooting of the EXPPCS (Experiment of the Physics of Colloids in Space) has been added to tomorrow's crew timeline in order to recover a very valuable experiment specimen.

╩ Walz performed monthly maintenance on the TVIS treadmill, also checking for any mechanical binding or noise during manual belt rotation, and the crew used it for today's physical exercise. Carl Walz was scheduled for the IWIS data collection during the running exercise.

CDR Onufrienko spent about an hour cleaning SM thermal control system fan grids. He then worked on the SM Wiener laptop preparing log files for downlink, copying them from hard disk to diskette, which then was transferred to the OCA computer set up for file downlink to the ground.

Dan ended the SSAS/FMK (solid sorbent air sampler/formaldehyde monitoring kit) operation started 24 hours ago and performed data collection to the MEC after deactivating the equipment.

The two flight engineers completed the second part of the EPO (educational payload operations) experiment by performing four demos involving the application of tools in microgravity, using selected tools from PERS (payload equipment restraint system) and the IVA tool kit. The demos were recorded on video tape for educational use on the ground. The EPO camcorder setup was then disassembled.

Besides performing the regular daily routine maintenance tasks, all crewmembers did their weekly NTXN data collection on the encrypted "Interactions" mood, group and journal questionnaire on the HRF PC.

Yuri performed the periodic inspections in the U.S. segment (USOS), beginning with hatch seals and continuing with the PFEs (portable fire extinguishers) and PBAs (portable breathing apparatus) in Lab, Node and Airlock (A/L), as well as the ELPS (emergency lighting power source) in the Lab.

Avionics/software specialists on the ground are continuing their review of the MDM transition event yesterday to assure themselves that all systems responded appropriately and the events are completely understood.

EVA specialists are investigating the Metox (metal oxide) canister regeneration problem, having singled out five distinct issues requiring resolution/clarification: (a) could Metox act as catalyst in creating toxins from the air? (b) could a Metox canister absorb atmospheric constituents and then release them inside the EMU spacesuit? (c) could the charcoal inside the canisters ignite? (d) could the constituents collected on the charcoal ignite?, and (e) could MetoxÔ performance of Metox be reduced due to these chemicals?╩ A plan has been developed, and the possible 600-odd constituents narrowed down to a priority-ordered matrix for testing, estimated to take about two weeks. Experts are confident that Metox will be exonerated, but a backup plan based on LiOH (lithium hydroxide) cans will be in place for 8A to be safe.

The protocol for single-vent water dumps from the Lab in retrograde direction has now been signed by both sides, US and Russia (no such dump is planned for the rest of this year). A similar joint agreement for double-vent dumping will be worked next.

The Robotics and LCA (launch cradle assembly) inspection preparatory for 8A, postponed f">Pr2/27, has now been scheduled for next Thursday, 3/7.

The Vozdukh CO2 removal unit in the SM again exhibited a failure, which was encountered several times before. TsUP reloaded and restarted the system, which is now running again in Manual mode 5. Assistance by the US CDRA (carbon dioxide removal assembly) in the Lab was an option but not required.

TsUP/Moscow requests return on 8A of a sample of a "substance" detected by the crew on SM surfaces. Necessary paper work and travel arrangements for delivering the samples to Moscow are being made at MCC-H.

Two reboosts by Progress 6P +X thrusters are still scheduled for 3/6 and 3/13, each consisting of two burns (during crew sleep time, i.e., not requiring their assistance). Propellant from FGB will be used, transferred through feed lines going from FGB through the SM to the Progress at SM aft end. This also means that these prop lines will have to be purged and vented to space prior to undocking of 6P. The vent is scheduled for 3/15, and TsUP has requested use of the CSA SSRMS/Canadarm2 for video-viewing the event. The period between this date and the actual undocking on 3/19 will be used for setting up the "Kolibri" microsatellite experiment, checking pressurization, etc. Launch of 7P will then follow on 3/22.

Today's CEO (crew earth observation) target areas were Tigris-Euphrates, Turkey (of interest: oblique, detailed context views of this headwaters region, especially to document the mountain snow pack indicated in satellite imagery to the right of track), E. Mediterranean Dust and Smog (crew was asked to look for aerosol buildup in the weak high-pressure area situated over the eastern Mediterranean this pass, and to try for oblique views to the left of track towards the Balkan Peninsula and the isle of Crete), Gulf of St. Lawrence (clearing from the SW was expected by the time of this pass. Of interest: oblique context views to the right of track to document the extent of pack ice as the latest outbreak of Arctic air moves into this region), Eastern United States (as ISS crossed the lower peninsula of Michigan from the northwest, crew was to try for oblique views to either side of track looking or aerosols and documenting the fresh terrain and lake-effect snowfall patterns in the wake of yesterday's winter storm).

U.S. and Russian Segment Status (as of 2:01 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16 Amp mode, the lowest possible setting). Vozdukh CO2 scrubber is in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.
  • BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, #2 in Purify mode.
  • SM Working Compartment: Pressure (mmHg) -- 760, temperature (deg C) -- 26.5, ppO2 (mmHg) -- 154.5, ppCO2 (mmHg) -- 4.4.
  • SM Transfer Compartment: Pressure (mmHg) -- 762, temperature (deg C) -- 21.0; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • FGB Cabin: Pressure (mmHg) -- 756, temperature (deg C) -- 24.0; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Node: Pressure (mmHg) -- 753.43, temperature (deg C) -- 23.8 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab: Pressure (mmHg) -- 755.52, temperature (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a;
  • Joint Airlock (Equip. Lock): Pressure (mmHg) -- 755.32, temperature (deg C) -- n/a; shell heater temp (deg C) -- 21.3, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • PMA-1: Shell heater temp (deg C) -- 23.5.
  • PMA-2: Shell heater temp (deg C) -- 13.5.
(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data].

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Autotrack mode (during XPOP).
  • SM batteries: battery #3 in "cycle" mode, all others (7) in "partial charge" mode.
  • FGB battery #5 is offline; all other batteries (5) are in "partial charge" mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby, all with new R2 s/w.
  • GNC-2 MDM is prime; GNC-1 is back-up, with R1 s/w.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.
  • Ku-band is operating nominally in open loop pointing mode.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) operating nominally.

Robotics:

  • SSRMS/Canadarm2 at Port Stow position, with Keep Alive power on both strings.
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 6:15am EST):

  • Mean altitude -- 383.1 km
  • Apogee -- 390.2 km
  • Perigee -- 376.0 km
  • Period -- 92.2 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.0010505
  • Orbits per 24-hr. day -- 15.62
  • Altitude decrease -- 304 m (mean) in last 19 hours.
  • Solar Beta Angle: -57.4 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. '98) -- 18677
  • Current Flight Attitude -- XPOP (x-axis perpendicular to orbit plane [yaw: ~0 deg, pitch: -5 deg., roll: 0 deg])
For more on ISS orbit and naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

Please follow SpaceRef on Twitter and Like us on Facebook.